OpenStudy (anonymous):

Find the indefinite integral of (5cosx-2sec^2x)dx

OpenStudy (amistre64):

who lost it?

OpenStudy (amistre64):

its a sum so we could just find the int of each term

OpenStudy (amistre64):

sec^2 might be tricky tho

OpenStudy (amistre64):

tan goes to sec^2 tho so not all that tricky

OpenStudy (mr.math):

Does "goes to" mean "is the integral of"?

OpenStudy (amistre64):

derivatives are down to me; and ints are up; tan goes down to sec^2 in my mind

OpenStudy (mr.math):

Hmm interesting :-D

OpenStudy (amistre64):

might be the power thing; x^3 does down a degree ...

OpenStudy (amistre64):

5cos goes up to 5sin -2sec^2 goes up to -2tan then we tack on the +C for indefininity

OpenStudy (anonymous):

I can't get it... :/ I know the integral of cosine is sinx+c and the derivative of sec^2x is tanx+c

OpenStudy (anonymous):

*integral

OpenStudy (amistre64):

then thats it; the +Cs just merge into some uber constant

OpenStudy (amistre64):

\[\int (5cos(x)-2sec^2(x))dx\] \[\int (5cos(x))dx-\int (2sec^2(x))dx\] \[5sin(x)+c_1-2tan(x)+c_2\] \[5sin(x)+-2tan(x)+C\]

OpenStudy (anonymous):

THANKS!

OpenStudy (amistre64):

yw